# Improving Early Detection of Gravitational Waves from Binary Neutron Stars Using CNNs and FPGAs

### v2

@@ IV-C Learning strategy
TABLE I: In this classification task the positive class is a GW simulated signal added in raw detector noise. To enhance the learning we create five curriculum learning data sets $D$. In this table, we present the most relevant magnitudes of the GW simulations for each curriculum learning data set: average minimum frequency $f_{min}$, maximum frequency $f_{max}$, network (Hanford, Livingston and Virgo) SNR and PISNR.
@@
Here, $\cup$ denotes the set union of the data, and the curriculum step is $k=\{1,2,3,4,5\}$. Each curriculum step in FindCNN lasts 6 epochs, minimizing a weighted BCEWithLogitsLoss to lower the number of FPs, i.e. it is more relevant to correctly classify the noise class (negative class) than the injection class (positive class) to avoid false alerts. Thus, as in the original work, we weigh the injection class by 0.4. Regarding the optimizer, we employ AdaMax [70] with a weight decay of $10^{-5}$ and the learning rate is $8\times 10^{-5}$.
@@
where $\mathcal{L}^{e}_{val}$ represents the validation loss of the current epoch $e$, $\mathcal{L}^{e^{\text{best}}}_{val}$ is the validation loss of the best epoch $e^{\text{best}}$, $\epsilon$ the tolerance and $n$ the patience. After several experiments we set $\epsilon=10^{-4}$ and $n=2$. Moreover, for each curriculum step, we select the weights of the best-performing epoch, i.e. the one with the lowest validation loss, as initialization for the next step. We also employ BCEWithLogitsLoss weighted at $0.1$. After several experiments, AdamW optimizer [70] showed the best performance. Furthermore, we use an adaptive learning rate to allow the network to learn more slowly at more difficult steps: given an initial learning rate $lr_{0}=\times 10^{5}$, at each curriculum step $c\in[1,5]$, we set $lr_{c}=lr_{0}/c$.
IV-D Modified architectures and quantization
In this work, we used the AMD Kria KV260 Vision AI Starter Kit [62], and the Vitis AI software [71] for the deployment of the models. As Kria KV260 is designed for vision applications, it is only prepared to deal with 2-dimensional operations. Hence, to run the models in the chosen FPGA, we transformed the 1-dimensional FindCNN and GWaveNet to 2-dimensional versions: FindCNN2D and GWaveNet2D. Nonetheless, in a low-latency context, this implies reshaping within the FPGA, but such operation is not supported to run on the data processing unit (DPU), needing to be run in the non-optimized CPU of the FPGA.
@@ IV-D Modified architectures and quantization
Last but not least, due to the limited memory of FPGAs, it is necessary to quantize our models, transforming the model’s weights from floating point to lower bit-widths to save memory and accelerate inference. In this work, we use the power-of-two quantization, which is a robust logarithm-based quantization strategy implemented in Vitis AI [72].
V Results
V-A Performance
In Section III, we mentioned that the input data has $155,648\,$data points and three channels, corresponding to Hanford, Livingston and Virgo, that we use to train, validate and test both models. In Fig. 3, we present the losses (solid line) and accuracies (dashed line) for FindCNN (top panel) and GWaveNet (bottom panel) for training (blue) and validation (yellow). We use a decision threshold of $0.5$ to calculate the accuracies, i.e. the cutoff value used to distinguish the classes is 0.5. The progressive curriculum learning step is shown by vertical dotted lines, where at each step we add samples further away from the merger. Furthermore, we mark the best-performing epoch from each curriculum learning step in red, i.e. the epoch with the best validation loss.
@@ V-A Performance
Such overfitting is not observed in GWaveNet, as training and validation are close together throughout the progressive curriculum learning steps. Additionally, the best-performing epoch typically occurs in the middle or at the end of each step. Note that, due to the early stopping method (see Section IV-C), the curriculum steps are shorter for easier-to-learn datasets, and longer for harder ones. Another notable detail is the expected increase in loss at the beginning of each curriculum learning step, followed by a decrease as learning stabilises, which is characteristic of curriculum learning strategies.
@@
In a GW experiment, it is crucial assess the number of FP. A common metric in GW is the false alarm probability (FAP), also known as false positive rate. In the top panel of Fig. 6 we plot FAP as a function of the decision threshold for FindCNN (green) and GWaveNet (pink). We mark the target FAP $=1\%$ in red, as defined in [26]. We can observe that GWaveNet achieves FAP $=1\%$ for a decision threshold of 0.46, while FindCNN needs a more aggressive decision threshold of 0.86. Furthermore, GWaveNet allows to set FAP $\leq 10^{-4}$.
@@
One of the the main limitations of FindCNN (see Section II), is that the network forgets the previous curriculum learning steps. To compare the “memory” of FindCNN, we reproduce Fig. 5 of [26] in Fig. 7. Here, we represent the true alarm probability (TAP), also known as the true positive rate, as a function of the average PISNR for the final models after the progressive curriculum learning strategy. In the top panel of Fig. 7, we can see that the final FindCNN has a TAP $>0.5$ for signals with PISNR $>30$ and $f_{max}={20,25,30}\,$Hz, the most difficult samples. As expected, TAP decreases for lower PISNR, which corresponds to the most quiet signals. Nonetheless, its performance drops $\sim 20$ percentage points (p.p.) for $f_{max}={35,40}\,$Hz at $\approx 25$ PISNR, due to catastrophic forgetting. In a realistic GW search, this would be unreliable, as the algorithm would not alert about the signals closer to the merger. On the other hand, we can observe how the final GWaveNet can “remember” easier data sets while maintaining a comparable performance to FindCNN for the most difficult data set $f_{max}=20\,$Hz.
As we present in Table II, GWaveNet outperforms FindCNN by 101010\,10p.p., but these models are unable to run on an FPGA due to memory limitations (see Section IV-D). In Table, II we also show the accuracies of the modified models before and after quantization. It is relevant to note that FindCNN2D and GWaveNet2D outperform their 1-dimensional versions by $\sim 3\,$p.p., which could imply that Pytorch [64] is better optimized for vision applications. This could also result in better numerical stability, leading to higher accuracies, but further investigation is needed. On the other hand, the modified versions perform worse than the original ones, due to a simplification in certain operations. This is particularly notorious for GWaveNet2DModified, which has $\sim 5\,$p.p. less accuracy than original GWaveNet, but its performance improves $\sim 1\,$p.p. after quantization, likely due to the smoothing of the learnt manifold. Nonetheless, the worst-performing GWaveNet still outperforms the best-performing FindCNN by $\sim 2\,$p.p.
@@
TABLE II: Test accuracies at FAP set at 1% before and after quantization. All models were tested on 22000 samples, except for the GWaveNet2D and GWaveNet2DModified post-quantization models, which were tested on 5000 samples, due to timeout constraints. We highlight in bold the best FindCNN model and the worst GWaveNet model.
@@
TABLE III: Average time (ms) and energy consumption (mJ) to predict a single sample and economic cost of running each model in a year (\$currency-dollar\\$\$/year) for different devices (see details in Section IV-A) at 3 standard deviations given 10 experiments. Note that only FindCNN2D and FindCNN2DModified was compatible with our quantization framework.
V-B Time, energy and economic cost
To compare the models’ efficiencies, we measured the time taken to test 22,000 samples on both the CPU and GPU and 1,000 samples on the FPGA. The results of time (ms), energy consumption (mJ), and cost (\$currency-dollar\\$\$/year) to run each model for CPU, GPU and FPGA are presented in Table III. As mentioned in Section IV-D, we were unable to compile GWaveNet in the FPGA due to software limitations, potentially due to the lack of support of the padding needed for causal convolutions, but further investigation is required.
Regarding the average inference time, we can observe how all FindCNN models take $\sim 25\,$ms to predict a sample in CPU and GPU. However, the time in an FPGA is 4 times longer. Further exploration revealed that the bottleneck was caused by the batch normalization layer in Vitis AI [73]. Ignoring this layer decreased the FPGA times and energy consumption in half. Nonetheless, the lack of speed of the FPGA could be due to our particular low-end FPGA, which was not designed for processing high-dimensional time series data.
@@ V-B Time, energy and economic cost
Regarding the energy consumption needed for inference (see Table III), when comparing the same model across various devices, we can see that while the GPU is one of the fastest, it is also the one that consumes the most energy. As before, if we ignore the batch normalization layers of FindCNN models on the FPGA, the energy consumption is halved, being comparable to the CPU consumption. Thus, we believe that a model fully optimized for an FPGA would have a better energy performance, and this will be explored in future works.
In a realistic situation, these models would run through the whole observing run. In Table III, we consider the economic cost of running the models for a year, taking the average industry electricity price in the USA, as most GW clusters are there. In 2022, the average electricity cost was 0.08450.08450.0845\,0.0845\$/kWh [74]. As expected, the GPU is the most expensive. On the other hand, the cost of running on an FPGA is $\approx 7\%$ cheaper than running on a CPU. Note that we only take into account the electricity prices, but adding the upfront and upkeep prices of the devices, the FPGA is by far the least expensive. Assuming the price at launch date over the number of lifetime years, under heavy use, and the prices for running FindCNN2D for a year, running on a CPU ($\approx 495\$$) or GPU ($\approx 465\$$) have similar expenses, whereas running on an FPGA ($\approx 65\$$) costs over seven times less.
@@
VI Conclusions
In this work to develop a novel ML-based GW detection algorithm for the early inspiral of BNS mergers, GWaveNet, inspired by the audio generating WaveNet [63]. We compare its performance to the current state-of-the-art, FindCNN [26], reproducing this work in simulated detector noise from the third observing run. Moreover, we compare their time, energy consumption and cost in CPU, GPU and FPGA.
As in [26], we implement a progressive curriculum learning strategy, where easier (more difficult) samples have higher (lower) maximum frequency $f_{max}$, and are closer (further) to the merger, meaning that they will be detected later (earlier). Both FindCNN and GWaveNet are binary classifiers, that differentiate between noise (negative class) and noise + signal (positive class) solely utilizing the inspiral of BNS mergers.

### v3

@@ 4.3 Learning strategy
Table 1: In this classification task the positive class is a GW simulated signal added in raw detector noise. To enhance the learning we create five curriculum learning data sets $D$. In this table, we present the most relevant magnitudes of the GW simulations for each curriculum learning data set: average minimum frequency $f_{min}$, maximum frequency $f_{max}$, network (Hanford, Livingston and Virgo) SNR and PISNR.
@@
Here, $\cup$ denotes the set union of the data, and the curriculum step is $k=\{1,2,3,4,5\}$. Each curriculum step in FindCNN lasts 6 epochs, minimizing a weighted BCEWithLogitsLoss to lower the number of FPs, i.e. it is more relevant to correctly classify the noise class (negative class) than the injection class (positive class) to avoid false alerts. Thus, as in the original work, we weigh the injection class by 0.4. Regarding the optimizer, we employ AdaMax Kingma and Ba (2017) with a weight decay of $10^{-5}$ and the learning rate is $8\times 10^{-5}$.
@@
where $\mathcal{L}^{e}_{val}$ represents the validation loss of the current epoch $e$, $\mathcal{L}^{e^{\text{best}}}_{val}$ is the validation loss of the best epoch $e^{\text{best}}$, $\epsilon$ the tolerance and $n$ the patience. After several experiments we set $\epsilon=10^{-4}$ and $n=2$. Moreover, for each curriculum step, we select the weights of the best-performing epoch, i.e. the one with the lowest validation loss, as initialization for the next step. We also employ BCEWithLogitsLoss weighted at $0.1$. After several experiments, AdamW optimizer Kingma and Ba (2017) showed the best performance. Furthermore, we use an adaptive learning rate to allow the network to learn more slowly at more difficult steps: given an initial learning rate $lr_{0}=\times 10^{5}$, at each curriculum step $c\in[1,5]$, we set $lr_{c}=lr_{0}/c$.
4.4 Modified architectures and quantization
In this work, we used the AMD Kria KV260 Vision AI Starter Kit Advanced Micro Devices, Inc. (Accessed 27 March 2024), and the Vitis AI software Advanced Micro Devices, Inc. (Accessed 02 April 2024) for the deployment of the models. As Kria KV260 is designed for vision applications, it is only prepared to deal with 2-dimensional operations. Hence, to run the models in the chosen FPGA, we transformed the 1-dimensional FindCNN and GWaveNet to 2-dimensional versions: FindCNN2D and GWaveNet2D. Nonetheless, in a low-latency context, this implies reshaping within the FPGA, but such operation is not supported to run on the data processing unit (DPU), needing to be run in the non-optimized CPU of the FPGA.
@@ 4.4 Modified architectures and quantization
Last but not least, due to the limited memory of FPGAs, it is necessary to quantize our models, transforming the model’s weights from floating point to lower bit-widths to save memory and accelerate inference. In this work, we use the power-of-two quantization, which is a robust logarithm-based quantization strategy implemented in Vitis AI Xilinx, Inc. (Accessed 07 May 2024).
5 Results
5.1 Performance
In Section 3, we mentioned that the input data has $155,648\,$data points and three channels, corresponding to Hanford, Livingston and Virgo, that we use to train, validate and test both models. In Fig. 3, we present the losses (solid line) and accuracies (dashed line) for FindCNN (top panel) and GWaveNet (bottom panel) for training (blue) and validation (yellow). We use a decision threshold of $0.5$ to calculate the accuracies, i.e. the cutoff value used to distinguish the classes is 0.5. The progressive curriculum learning step is shown by vertical dotted lines, where at each step we add samples further away from the merger. Furthermore, we mark the best-performing epoch from each curriculum learning step in red, i.e. the epoch with the best validation loss.
@@ 5.1 Performance
Such overfitting is not observed in GWaveNet, as training and validation are close together throughout the progressive curriculum learning steps. Additionally, the best-performing epoch typically occurs in the middle or at the end of each step. Note that, due to the early stopping method (see Section 4.3), the curriculum steps are shorter for easier-to-learn datasets, and longer for harder ones. Another notable detail is the expected increase in loss at the beginning of each curriculum learning step, followed by a decrease as learning stabilises, which is characteristic of curriculum learning strategies.
@@
In a GW experiment, it is crucial assess the number of FP. A common metric in GW is the false alarm probability (FAP), also known as false positive rate. In the top panel of Fig. 6 we plot FAP as a function of the decision threshold for FindCNN (green) and GWaveNet (pink). We mark the target FAP $=1\%$ in red, as defined in Baltus et al. (2022). We can observe that GWaveNet achieves FAP $=1\%$ for a decision threshold of 0.46, while FindCNN needs a more aggressive decision threshold of 0.86. Furthermore, GWaveNet allows to set FAP $\leq 10^{-4}$.
@@
One of the the main limitations of FindCNN (see Section 2), is that the network forgets the previous curriculum learning steps. To compare the “memory” of FindCNN, we reproduce Fig. 5 of Baltus et al. (2022) in Fig. 7. Here, we represent the true alarm probability (TAP), also known as the true positive rate, as a function of the average PISNR for the final models after the progressive curriculum learning strategy. In the top panel of Fig. 7, we can see that the final FindCNN has a TAP $>0.5$ for signals with PISNR $>30$ and $f_{max}={20,25,30}\,$Hz, the most difficult samples. As expected, TAP decreases for lower PISNR, which corresponds to the most quiet signals. Nonetheless, its performance drops $\sim 20$ percentage points (p.p.) for $f_{max}={35,40}\,$Hz at $\approx 25$ PISNR, due to catastrophic forgetting. In a realistic GW search, this would be unreliable, as the algorithm would not alert about the signals closer to the merger. On the other hand, we can observe how the final GWaveNet can “remember” easier data sets while maintaining a comparable performance to FindCNN for the most difficult data set $f_{max}=20\,$Hz.
As we present in Table 2, GWaveNet outperforms FindCNN by 101010\,10p.p., but these models are unable to run on an FPGA due to memory limitations (see Section 4.4). In Table, 2 we also show the accuracies of the modified models before and after quantization. It is relevant to note that FindCNN2D and GWaveNet2D outperform their 1-dimensional versions by $\sim 3\,$p.p., which could imply that Pytorch A. et al. (2019) is better optimized for vision applications. This could also result in better numerical stability, leading to higher accuracies, but further investigation is needed. On the other hand, the modified versions perform worse than the original ones, due to a simplification in certain operations. This is particularly notorious for GWaveNet2DModified, which has $\sim 5\,$p.p. less accuracy than original GWaveNet, but its performance improves $\sim 1\,$p.p. after quantization, likely due to the smoothing of the learnt manifold. Nonetheless, the worst-performing GWaveNet still outperforms the best-performing FindCNN by $\sim 2\,$p.p.
@@
Table 2: Test accuracies at FAP set at 1% before and after quantization. All models were tested on 22000 samples, except for the GWaveNet2D and GWaveNet2DModified post-quantization models, which were tested on 5000 samples, due to timeout constraints. We highlight in bold the best FindCNN model and the worst GWaveNet model.
@@
Table 3: Average time (ms) and energy consumption (mJ) to predict a single sample and economic cost of running each model in a year (\$currency-dollar\\$\$/year) for different devices (see details in Section 4.1) at 3 standard deviations given 10 experiments. Note that only FindCNN2D and FindCNN2DModified was compatible with our quantization framework.
5.2 Time, energy and economic cost
To compare the models’ efficiencies, we measured the time taken to test 22,000 samples on both the CPU and GPU and 1,000 samples on the FPGA. The results of time (ms), energy consumption (mJ), and cost (\$currency-dollar\\$\$/year) to run each model for CPU, GPU and FPGA are presented in Table 3. As mentioned in Section 4.4, we were unable to compile GWaveNet in the FPGA due to software limitations, potentially due to the lack of support of the padding needed for causal convolutions, but further investigation is required.
Regarding the average inference time, we can observe how all FindCNN models take $\sim 25\,$ms to predict a sample in CPU and GPU. However, the time in an FPGA is 4 times longer. Further exploration revealed that the bottleneck was caused by the batch normalization layer in Vitis AI Advanced Micro Devices, Inc. (Accessed 2024). Ignoring this layer decreased the FPGA times and energy consumption in half. Nonetheless, the lack of speed of the FPGA could be due to our particular low-end FPGA, which was not designed for processing high-dimensional time series data.
@@ 5.2 Time, energy and economic cost
Regarding the energy consumption needed for inference (see Table 3), when comparing the same model across various devices, we can see that while the GPU is one of the fastest, it is also the one that consumes the most energy. As before, if we ignore the batch normalization layers of FindCNN models on the FPGA, the energy consumption is halved, being comparable to the CPU consumption. Thus, we believe that a model fully optimized for an FPGA would have a better energy performance, and this will be explored in future works.
In a realistic situation, these models would run through the whole observing run. In Table 3, we consider the economic cost of running the models for a year, taking the average industry electricity price in the USA, as most GW clusters are there. In 2022, the average electricity cost was 0.08450.08450.0845\,0.0845\$/kWh Statista (2024). As expected, the GPU is the most expensive. On the other hand, the cost of running on an FPGA is $\approx 7\%$ cheaper than running on a CPU. Note that we only take into account the electricity prices, but adding the upfront and upkeep prices of the devices, the FPGA is by far the least expensive. Assuming the price at launch date over the number of lifetime years, under heavy use, and the prices for running FindCNN2D for a year, running on a CPU ($\approx 495\$$) or GPU ($\approx 465\$$) have similar expenses, whereas running on an FPGA ($\approx 65\$$) costs over seven times less.
@@
6 Conclusions
In this work to develop a novel ML-based GW detection algorithm for the early inspiral of BNS mergers, GWaveNet, inspired by the audio generating WaveNet van den Oord et al. (2016c). We compare its performance to the current state-of-the-art, FindCNN Baltus et al. (2022), reproducing this work in simulated detector noise from the third observing run. Moreover, we compare their time, energy consumption and cost in CPU, GPU and FPGA.
As in Baltus et al. (2022), we implement a progressive curriculum learning strategy, where easier (more difficult) samples have higher (lower) maximum frequency $f_{max}$, and are closer (further) to the merger, meaning that they will be detected later (earlier). Both FindCNN and GWaveNet are binary classifiers, that differentiate between noise (negative class) and noise + signal (positive class) solely utilizing the inspiral of BNS mergers.